OpenStudy (anonymous):

Help :) (3+ radical 5) (2- radical 2)

OpenStudy (anonymous):

Basic FOIL: \[ (3 + \sqrt{5})(2 - \sqrt{2}) = \boxed {6 + 2\sqrt{5} - 3\sqrt{2} - \sqrt{10}}. \]

OpenStudy (anonymous):

3.067 or did were you asking the method of foiling

OpenStudy (anonymous):

ok thank youuu

OpenStudy (anonymous):

you do 3 times 2, then 3 times (-rad2), then rad5 times 2, and finally rad5 times -rad2

OpenStudy (anonymous):

ohh alright thanks i get it now :)